# Beyond Autoregression: Fast LLMs via Self-Distillation Through Time

### v1

@@ Abstract
Autoregressive (AR) Large Language Models (LLMs) have demonstrated significant success across numerous tasks. However, the AR modeling paradigm presents certain limitations; for instance, contemporary autoregressive LLMs are trained to generate one token at a time, which can result in noticeable latency. Recent advances have indicated that search and repeated sampling can enhance performance in various applications, such as theorem proving, code generation, and alignment, by utilizing greater computational resources during inference. In this study, we demonstrate that diffusion language models are capable of generating at least 32 tokens simultaneously, while exceeding the performance of AR models in text quality and on the LAMBADA natural language understanding benchmark. This outcome is achieved through a novel distillation method for discrete diffusion models, which reduces the number of inference steps by a factor of 32-64. Practically, our models, even without caching, can generate tokens at a rate that is up to 8 times faster than AR models employing KV caching, and we anticipate further improvements with the inclusion of caching. Moreover, we demonstrate the efficacy of our approach for diffusion language models with up to 860M parameters.
@@ 1 Introduction
Importantly, the benefits of repeated sampling can be so significant that it is often more efficient to use a smaller, faster model rather than a larger, slower one. More generally, one can improve the performance of a fixed model by scaling up computational resources at inference time (Madaan et al., 2023; Yao et al., 2023; Snell et al., 2024; Wu et al., 2024; Chen et al., 2024; Brown et al., 2024; Goyal et al., 2024), a phenomenon that was previously observed for games (Campbell et al., 2002; Silver et al., 2016; Lerer et al., 2019; Brown et al., 2020a; Jones, 2021). Hence, when tackling reasoning tasks, a major bottleneck is the latency of the model. In this work, we improve the decoding speed of LLMs by moving away from AR modeling. We build on recent breakthroughs in discrete diffusion (Lou et al., 2023; Sahoo et al., 2024; Shi et al., 2024; Ou et al., 2024). Our approach can generate text up to 8 times faster than AR models that use KV caching (Pope et al., 2022). Diffusion models are typically trained to maximize the evidence lower bound (ELBO), which does not consider the desired number of inference steps. To improve text quality with fewer forward passes, we train a student model to match the distribution of samples produced using a large number of steps. This is a significant finding because faster decoding mechanisms allow exploring a larger search space in applications that require search, planning, and reranking. In summary, our core contributions are as follows:
  • We introduce Self-Distillation Through Time (SDTT), which allows generating at least 32 tokens at a time, while achieving better perplexity than GPT-2 with nucleus sampling. Unlike many distillation methods for continuous diffusion models, SDTT does not rely on deterministic mappings such as DDIM (Song et al., 2022). SDTT is very simple and easy to implement.
  • We show that SDTT can generate tokens up to 8 times faster than AR models that use KV caching. In particular, our model does not rely on activation caching, suggesting that there is potential for even greater efficiency gains.
  • We demonstrate the effectiveness of SDTT for language models with up to 860M parameters. To the best of our knowledge, this is the largest publicly available discrete diffusion language model.
@@ 2 Background
Figure 1: Performance on LAMBADA after multiple rounds of SDTT with different distillation losses. We pre-train with the masked diffusion language modeling objective (MDLM) (Sahoo et al., 2024) and distill with 7 rounds of SDTT. The sampling step size (x-axis) denotes the expected number of decoding steps to match the MDLM teacher. For example, 1/256 denotes the student after 2 rounds of SDTT. Note that a single word in the LAMBADA data set often consists of multiple tokens.
@@ 2.1 Masked diffusion language modeling
We follow the notation of Sahoo et al. (2024) to introduce masked diffusion language modeling (MDLM.) Language modeling can be framed as the sequential prediction task of discrete tokens ($x_{i}$) coming from a vocabulary $\mathcal{X}=\mathbb{Z}^{<N}=\left\{0,~{}...,~{}N-1\right\}$ that can take $N$ possible discrete values. A language model would predict sequences of length $L$, which can be defined as the sequences of $x_{i}$’s originating from $\mathcal{X}^{L}=\left\{\mathbf{x}^{(i)}=(x^{(i)}_{0},~{}\dots,~{}x^{(i)}_{L-1})\right\}_{i\in\mathbb{Z}^{<K}}$. Let $\mathcal{D}:=\left\{\mathbf{x}^{(0)},~{}\dots,~{}\mathbf{x}^{(K-1)}:\mathbf{x}^{(i)}\in\mathcal{X}^{L}\right\}$ denote the training set. The goal of language modeling is to sample from the unknown distribution $p_{0}:\mathcal{X}^{L}\rightarrow[0,1]$ that generated the samples in $\mathcal{D}$.
@@
To generate new samples, we can simulate the reverse process from eq. 2. Since the ground-truth sample ${\mathbf{x}}$ is unknown, Sahoo et al. (2024) learn an approximation ${\mathbf{x}}_{\theta}$ using a neural network with parameters $\theta$. Sahoo et al. (2024) then use ${\mathbf{x}}_{\theta}$ instead of ${\mathbf{x}}$ to simulate the reverse process. The sampling distribution is denoted as ${p_{\theta}({\mathbf{z}}_{s}|{\mathbf{z}}_{t}):=q({\mathbf{z}}_{s}|{\mathbf{z}}_{t},{\mathbf{x}}_{\theta}({\mathbf{z}}_{t},t)})$. Sahoo et al. (2024) optimizes $\theta$ using a continuous version of the negative evidence lower bound (ELBO) of Sohl-Dickstein et al. (2015a). Previous research has shown that continuous-time objectives optimize the data likelihood better (Kingma et al., 2023). Due to the definition of the absorbing diffusion process, the ELBO simplifies to a weighted cross-entropy loss between the ground-truth ${\mathbf{x}}$ and the model predictions ${\mathbf{x}}_{\theta}$:
@@
(b) Latency of the distilled models. Sampling with 16 steps achieves an 8x speedup against the AR baseline that uses KV caching.
Figure 2: SDTT. In figure (a), we illustrate how we prepare the distillation targets. In figure (b), we show the latency results with 16, 32, and 64 sampling steps.
@@ 3.1 Self-Distillation Through Time
As explained in section 2.1, training the discrete diffusion language model involves optimizing the ELBO over the training examples. Fewer steps typically lead to lower sample quality because the approximation of the reverse process is less accurate, as visible in the teacher curve in fig. 3.
To address the issue of low sample quality with fewer steps, we propose Self-Distillation Through Time (SDTT). SDTT tunes a pre-trained MDLM to allow decoding with significantly fewer steps. Interestingly, our final model decodes samples with lower generative perplexity in 32 steps than the teacher would with 1024 forward passes. In short, SDTT improves the sampling speed by distilling the inference time computation to sample multiple steps into the student.
@@
where $\delta$ a divergence measure between the student and the objectives $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}}))$. We consider the Kullback-Leibler divergence (KLD), Total Variation Distance (TVD), and Mean-Squared Error (MSE). See appendix B for details on those divergence measures.
Algorithm 2 Self-Distillation Through Time training loop.
@@
Following the terminology of knowledge distillation, we call the denoiser ${\mathbf{x}}_{\theta}$ used for many steps decoding from the teacher and the denoiser ${\mathbf{x}}_{\nu}$ used for a few steps decoding from the student. To train ${\mathbf{x}}_{\nu}$ to match the predictions of ${\mathbf{x}}_{\theta}$, we sample from the teacher for the $\nicefrac{{m}}{{k}}$ steps. Whenever a MASK token is denoised, we collect the log probabilities predicted by the teacher for this MASK token. These log-probabilities become the distillation targets $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$. Algorithm 1 outlines this process and fig. 2 presents it visually. While fig. 2 shows the distillation for two steps, the procedure can be extended to larger values of $\nicefrac{{m}}{{k}}$. The complete SDTT training loop is presented in algorithm 2.
@@ 4 Experiments
We distill MDLMs on the OpenWebText dataset (Gokaslan & Cohen, 2019) as it was used to train recent discrete diffusion language models (Lou et al., 2023; Sahoo et al., 2024). We use the Adam optimizer with a learning rate of $56e-$ 5, a batch size of 128 and no weight decay. We linearly increase the learning rate for 500 steps and keep it constant afterwards. As a base model, we reuse the checkpoint released by Sahoo et al. (2024). See appendix C for more details.
In section 4.1, we evaluate 3 distillation divergences and show that SDTT can be applied iteratively to reduce the number of sampling steps by a factor 16-32. In section 4.2, we ablate on the importance of hyperparameters, including the duration of each SDTT round and the number of sampling steps to generate $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$. In fig. 5, we scale SDTT to models with of up to 860M parameters. Finally, in section 4.4, we compare the latency of SDTT against autoregressive models that use KV caching.
Unconditional generation
Following prior work (Dieleman et al., 2022; Lou et al., 2023; Sahoo et al., 2024), we use a larger model to compute the perplexity of generated unconditional text examples. We evaluate the smallest students using GPT-2 (large) (Radford et al., 2019). In the scaling experiments, we use Llama3 8B (Touvron et al., 2023), since we compare models with up to 860M parameters. We decode 1024 samples of 1024 tokens using from 16 to 1024 sampling steps, and compare the perplexity of samples generated by an AR model trained from scratch, as well as GPT2 using nucleus sampling with $p=0.95$.
Conditional generation
To evaluate conditional generation, we use the MAUVE score(Pillutla et al., 2021). MAUVE measures how well a model follows a prompt by comparing multiple generations with a reference continuation. We use the first 1024 samples with at least 1024 tokens from the WebText dataset (OpenAI, 2019), use the first 50 tokens as a prompt, and generate 50 tokens of continuation. For each prompt, we generate five continuations, as done in Lou et al. (2023).
@@
We use the LAMBADA dataset (Paperno et al., 2016) as a downstream evaluation benchmark. We report an upper bound on the perplexity, computed using the ELBO (3). The LAMBADA accuracy is computed by masking all tokens of the last word and predicting all of them in a single forward pass, using the argmax of the predictions. The diffusion model is correct only if all the masked tokens are decoded correctly in a single step.
(a) SDTT on small models trained for 1M steps. dt=1/k means students matching teacher samples in $k$ steps. Successive lines correspond to additional SDTT rounds. SDTT can outperform the teacher and GPT-2 with nucleus sampling.
(b) Scaling SDTT to 860M parameters. Each scale is distilled for 7 rounds and the performance of the final student is displayed. In all 3 experiments, the student and teacher have the same number of parameters. The small distilled student reaches lower perplexity than the large teacher.
@@
4.1 Training objective and Iterated SDTT
SDTT requires choosing a divergence $\delta$ and we study the Mean-Squared Error (MSE), Total Variation Distance (TVD) and (reverse) Kullback-Leibler Divergence (KLD). To further reduce the number of sampling steps, we apply SDTT multiple times, using the distilled student as the teacher for subsequent rounds of SDTT. We distill the original models for 7 rounds of $10k$ training steps, and generate $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$ with $2$ teacher steps (algorithm 1). In this experiment, we used an exponential moving average (EMA) of the weights with a decay of $0.9999$ that we do not reset between rounds.
Both Figure 1 and Figure 5 show that students distilled with the KLD clearly outperform students trained using the MSE and TVD. The LAMBADA accuracy of students tuned with the KLD slightly improves over the teacher, while the perplexity remains better or matches the AR baselines for all but the last round of SDTT. The improved accuracy on LAMBADA suggests that the model is better at predicting multiple tokens in parallel after distillation with SDTT, since we evaluates the accuracy by decoding all tokens of the last word simultaneously. Figure 4 shows that the KLD seem to outperform the MSE and TVD objectives on MAUVE. Since we generate sequences of 100 tokens only for MAUVE, following (Lou et al., 2023), we sample with at most 128 steps, and use samples generated with 128 sampling steps from the teacher as a baseline. Note that as observed by Deschenaux & Gulcehre (2024), discrete diffusion models typically achieve slightly lower MAUVE scores than AR models. Nonetheless, distillation with the KLD objective improves the MAUVE score of the students. Figure 5 shows that students distilled with KLD have higher generative perplexity than those distilled with the MSE. However, KLD is the only objective that preserves performance in the LAMBADA data set while still significantly reducing the generative perplexity compared to the teacher. Therefore, in the remainder of this work, we focus on the KLD.

### v2

@@ Abstract
Autoregressive (AR) Large Language Models (LLMs) have demonstrated significant success across numerous tasks. However, the AR modeling paradigm presents certain limitations; for instance, contemporary autoregressive LLMs are trained to generate one token at a time, which can result in noticeable latency. Recent advances have indicated that search and repeated sampling can enhance performance in various applications, such as theorem proving, code generation, and alignment, by utilizing greater computational resources during inference. In this study, we demonstrate that diffusion language models are capable of generating at least 32 tokens simultaneously, while exceeding the performance of AR models in text quality and on the LAMBADA natural language understanding benchmark. This outcome is achieved through a novel distillation method for discrete diffusion models, which reduces the number of inference steps by a factor of 32-64. Practically, at the 1.3B parameters scale, diffusion models, even without caching, can generate tokens at a rate that is up to 8 times faster than AR models employing KV-caching, and we anticipate further improvements with the inclusion of caching. Moreover, we demonstrate the efficacy of our approach for diffusion language models with up to 860M parameters.
@@ 1 Introduction
Figure 1: Perplexity versus latency. The diffusion models (169M) use 16, 32, 64, 128 and 256 decoding step.
@@
Importantly, the benefits of repeated sampling can be so significant that it is often more efficient to use a smaller, faster model rather than a larger, slower one. More generally, one can improve the performance of a fixed model by scaling up computational resources at inference time (Madaan et al., 2023; Yao et al., 2023; Snell et al., 2024; Wu et al., 2024; Chen et al., 2024; Brown et al., 2024; Goyal et al., 2024), a phenomenon that was previously observed for games (Campbell et al., 2002; Silver et al., 2016; Lerer et al., 2019; Brown et al., 2020a; Jones, 2021). Hence, when tackling reasoning tasks, a major bottleneck is the latency of the model. In this work, we improve the decoding speed of LLMs by moving away from AR modeling. We build on recent breakthroughs in discrete diffusion (Lou et al., 2023; Sahoo et al., 2024; Shi et al., 2024; Ou et al., 2024). Our approach can generate text up to 8 times faster than AR models that use KV caching (Pope et al., 2022). Diffusion models are typically trained to maximize the evidence lower bound (ELBO), which does not consider the desired number of inference steps. Hence, vanilla diffusion models typically require thousands of decoding steps. Fortunately, it is possible to drastically reduce the inference costs of continuous diffusion models via distillation (Luhman & Luhman, 2021; Salimans & Ho, 2022). Continuous distillation methods rely on deterministic mappings from noise to data, such as DDIM (Song et al., 2022). The deterministic mappings can be efficiently learned by a student diffusion model to sample in fewer steps. We hypothesize that such deterministic map cannot exist for the diffusion language models studied in this work. Indeed, those models always initialize the denoising process with a sequence of masked token, hence a deterministic algorithm can only generate a single sample. As such, we devise a distillation method that does not does depend on deterministic maps. This is a significant finding because faster decoding mechanisms allow exploring a larger search space in applications that require search, planning, and reranking. In summary, our core contributions are as follows:
  • We introduce Self-Distillation Through Time (SDTT), which allows generating at least 32 tokens at a time, while achieving better perplexity than GPT-2 with nucleus sampling for conditional and unconditional generation. Unlike many distillation methods for continuous diffusion models, SDTT does not rely on deterministic mappings such as DDIM (Song et al., 2022). SDTT is very simple and easy to implement.
  • We show that SDTT can generate tokens up to 8 times faster than AR models that use KV caching, for models with 1.3B parameters, in 16 decoding steps. Importantly, the discrete diffusion model does not rely on activation caching, suggesting that there is potential for even greater efficiency gains. The latency gains for smaller models are even greater.
  • We demonstrate the effectiveness of SDTT for models with up to 860M parameters. To the best of our knowledge, this represents the largest publicly available discrete diffusion language model.
  • We evaluate the distilled students on LAMBADA (Paperno et al., 2016) and 6 multiple-choice questions benchmarks from Gao et al. (2021). We find that SDTT preserves the natural language understanding performance of the teacher.
@@ 2 Background
Figure 2: Performance on LAMBADA after multiple rounds of SDTT with different distillation losses. We pre-train with the masked diffusion language modeling objective (MDLM) (Sahoo et al., 2024) and distill with 7 rounds of SDTT. Note that a single word in the LAMBADA data set often consists of multiple tokens. We greedily decode all tokens a single forward pass for the diffusion models and decode autoregressively for the AR models.
@@ 2.1 Masked diffusion language modeling
(b) SDTT on small models trained for 1M steps. Successive lines correspond to additional SDTT rounds. SDTT can outperform the teacher and GPT-2 with nucleus sampling.
Figure 3: SDTT. In figure (a), we illustrate how we prepare the distillation targets. In figure (b), we display the generative perplexity of samples after distillation.
We follow the notation of Sahoo et al. (2024) to introduce masked diffusion language modeling (MDLM). Language modeling can be framed as the sequential prediction task of discrete tokens ($x_{i}$) coming from a vocabulary $\mathcal{X}=\mathbb{Z}^{<N}=\left\{0,~{}...,~{}N-1\right\}$ that can take $N$ possible discrete values. A language model would predict sequences of length $L$, which can be defined as the sequences of $x_{i}$’s originating from $\mathcal{X}^{L}=\left\{\mathbf{x}^{(i)}=(x^{(i)}_{0},~{}\dots,~{}x^{(i)}_{L-1})\right\}_{i\in\mathbb{Z}^{<K}}$. Let $\mathcal{D}:=\left\{\mathbf{x}^{(0)},~{}\dots,~{}\mathbf{x}^{(K-1)}:\mathbf{x}^{(i)}\in\mathcal{X}^{L}\right\}$ denote the training set. The goal of language modeling is to sample from the unknown distribution $p_{0}:\mathcal{X}^{L}\rightarrow[0,1]$ that generated the samples in $\mathcal{D}$.
@@
To generate new samples, we can simulate the reverse process from eq. 2. Since the ground-truth sample ${\mathbf{x}}$ is unknown, Sahoo et al. (2024) learn an approximation ${\mathbf{x}}_{\theta}$ using a neural network with parameters $\theta$. Sahoo et al. (2024) then use ${\mathbf{x}}_{\theta}$ instead of ${\mathbf{x}}$ to simulate the reverse process. The sampling distribution is denoted as ${p_{\theta}({\mathbf{z}}_{s}|{\mathbf{z}}_{t}):=q({\mathbf{z}}_{s}|{\mathbf{z}}_{t},{\mathbf{x}}_{\theta}({\mathbf{z}}_{t},t)})$. Sahoo et al. (2024) optimize $\theta$ using a continuous version of the negative evidence lower bound (NELBO) of Sohl-Dickstein et al. (2015a). Previous research has shown that continuous-time objectives optimize the data likelihood better (Kingma et al., 2023). Due to the definition of the absorbing diffusion process, the NELBO simplifies to a weighted cross-entropy loss between the ground-truth ${\mathbf{x}}$ and the model predictions ${\mathbf{x}}_{\theta}$:
@@ 3.1 Self-Distillation Through Time
As explained in fig. 3, discrete diffusion language models optimize the NELBO over the training examples. Fewer decoding steps typically lead to lower sample quality because the approximation of the reverse process is less accurate, as visible in the teacher curve in fig. 4.
To address the issue of low sample quality with fewer decoding steps, we propose Self-Distillation Through Time (SDTT). SDTT fine-tunes a pre-trained MDLM to allow decoding with significantly fewer steps. Interestingly, our final model decodes samples with lower generative perplexity in 32 steps than the teacher would with 1024 forward passes. In short, SDTT improves the sampling speed by distilling the inference time computation to sample multiple steps into the student.
@@
where $\delta$ a divergence measure between the student and the teacher targets $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}}))$. We consider the Kullback-Leibler divergence (KLD), Total Variation Distance (TVD), and Mean-Squared Error (MSE). See appendix B for details on those divergence measures.
Algorithm 2 One training round of Self-Distillation Through Time
@@
Following the terminology of knowledge distillation, we call the denoiser ${\mathbf{x}}_{\theta}$ used for many steps decoding as the teacher and the denoiser ${\mathbf{x}}_{\nu}$ used for a few steps decoding as the student. To train ${\mathbf{x}}_{\nu}$ to match the predictions of ${\mathbf{x}}_{\theta}$, we sample from the teacher for $\nicefrac{{m}}{{k}}$ steps. Whenever a MASK token is denoised, we collect the log probabilities predicted by the teacher for this MASK token. These log-probabilities become the distillation targets $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$. Algorithm 1 outlines this process and fig. 3(a) presents it visually. While fig. 3(a) shows how to distill two decoding steps in one, the procedure can be extended to larger values of $\nicefrac{{m}}{{k}}$. The complete SDTT training loop is presented in algorithm 2.
Iterated SDTT
SDTT reduces the number of decoding steps by a factor $\nicefrac{{m}}{{k}}$. If we want to reduce the number of decoding steps further, we can apply SDTT with $k^{\prime}<k$, or alternatively apply SDTT $n$ times, using the newly distilled student as teacher for the next round, which we refer to as iterated SDTT. Instead of directly optimizing the divergence in eq. 5, we introduce $n$ intermediate distributions $p_{\nu_{i}}^{k_{i}}$ such that $\nicefrac{{m}}{{k_{i}}}$ is an increasing sequence as a function of $i$. In practice, we choose $m=2^{10}$ and $k_{i}=2^{10-i}$ with $0\leq i\leq 7$ and sequentially minimize the objective
${\min_{\nu}~{}d\left(p_{\nu_{j}+1}^{(k_{j+1})}||p_{\nu_{j}}^{(k_{j})}\right),}$ (7)
for $70\leq j<$ 7, where $\nu_{j}$ denotes the parameters of the $j$-th denoiser, with $\nu_{0}=\theta$ (teacher). If the minimization procedure was perfect, minimizing eq. 5 or eq. 7 should result in the same solution. However in practice, we observe that it is easier to minimize eq. 7 sequentially for increasing values of $i$, in a progressive fashion, similar to Salimans & Ho (2022).
As an alternative to iterated SDTT, we tried using a single model and slowly growing the step size used to generate $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$. Unfortunately, this approach was unstable and the loss diverged after 30-50 steps, irrespective of how small the sampling step size was. Similar behavior was observed by Norouzi et al. (2023).
@@ 4 Experiments
We distill MDLMs on the OpenWebText dataset (Gokaslan & Cohen, 2019) as it was used to train recent discrete diffusion language models (Lou et al., 2023; Sahoo et al., 2024). We use the Adam optimizer with a learning rate of $56e-$ 5, a batch size of 128 and no weight decay. We linearly increase the learning rate for 500 training steps and keep it constant afterwards. As a base model, we reuse the checkpoint released by Sahoo et al. (2024). See appendix C for more details.
In section 4.1, we evaluate 3 distillation divergences and show that iterated SDTT can reduce the number of sampling steps by a factor 16-32. In section 4.2, we ablate on the importance of hyperparameters, including the duration of each round of iterated SDTT and the number of sampling steps to generate the targets $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$. In section 4.3, we scale SDTT to models with of up to 860M parameters. Finally, in section 4.4, we compare the latency of SDTT against autoregressive models that use KV caching.
Generative perplexity
Following prior work (Dieleman et al., 2022; Lou et al., 2023; Sahoo et al., 2024), we use a larger model to compute the generative perplexity of unconditional and conditional samples. We evaluate the smallest students using GPT-2 (large) (Radford et al., 2019). In the scaling experiments, we use Llama3 8B (Touvron et al., 2023), since we compare models with up to 860M parameters. As noted by Zheng et al. (2024a), the generative perplexity is sensitive to the floating-point precision. In this section, we sample using bfloat16, and report results using float64 in table 1. The conclusion are similar.
MAUVE
We evaluate conditional generation using the MAUVE score (Pillutla et al., 2021). MAUVE measures how well a model follows a prompt by comparing multiple generations with a reference continuation. We use the first 1024 samples with at least 1024 tokens from the WebText dataset (OpenAI, 2019), take the first 50 tokens as a prompt, and generate 50 tokens of continuation. For each prompt, we generate 5 continuations, as done in Lou et al. (2023).
Sample diversity
Post-training can drastically reduce the diversity of language models (Kirk et al., 2024; Agarwal et al., 2024; Li et al., 2024). Hence, we measure the diversity of samples using the self-BLEU score (Zhu et al., 2018) with the same completions used to compute MAUVE.
@@
We measure the downstream performance using the LAMBADA dataset (Paperno et al., 2016), as well as 6 multiple-choice question (MCQ) tasks from Gao et al. (2021). On LAMBADA, we report an upper bound on the perplexity, computed using the ELBO (3). We also report the suffix accuracy by masking all tokens of the last word and predicting all of them in a single forward pass, using the argmax of the predictions. The diffusion model is correct only if all the masked tokens are decoded correctly in a single decoding step. The 6 other benchmarks from Gao et al. (2021) evaluate the MCQ accuracy.
(a) Diversity of conditional generation (small scale). We measure the trade-off between quality and diversity using self-BLEU (Zhu et al., 2018). Deterministic sampling yields a score of 1. The diversity minimally decreases after distillation.
(b) Scaling SDTT to 860M parameters. The plot compares the performance of the teacher and final student (7 rounds). The student and teacher have the same size. The small distilled student reaches lower perplexity than the large teacher.
@@
4.1 Ablation on the training divergence
SDTT requires choosing a divergence $\delta$ and we study the Mean-Squared Error (MSE), Total Variation Distance (TVD) and (reverse) Kullback-Leibler Divergence (KLD). We apply iterated SDTT for 7 rounds of $10k$ training iterations and generate $\tilde{\mathbf{x}}_{\theta}^{\text{teacher}}({\mathbf{z}}_{t},t,\nicefrac{{m}}{{k}})$ with $2$ sampling steps from the teacher (algorithm 1). We use an exponential moving average (EMA) of the weights with a decay of $0.9999$ that we do not reset between rounds.
Figure 2 shows that students distilled with the KLD clearly outperform students trained using the MSE and TVD on LAMBADA. The LAMBADA accuracy of students tuned with the KLD slightly improves over the teacher, while the perplexity remains better or matches the AR baselines for all but the last round of SDTT. The improved accuracy on LAMBADA suggests that the model is better at predicting multiple tokens in parallel after distillation with SDTT, since we evaluates the accuracy by decoding all tokens of the last word simultaneously.
Figure 5 shows that the KLD seem to outperform the MSE and TVD objectives on MAUVE. Since we generate sequences of 100 tokens only for MAUVE, following (Lou et al., 2023), we sample with at most 128 steps, and use samples generated with 128 sampling steps from the teacher as a baseline. Note that as observed by Deschenaux & Gulcehre (2024), discrete diffusion models typically achieve slightly lower MAUVE scores than AR models. Nonetheless, distillation with the KLD objective improves the MAUVE score of the students. Similarly fig. 18 shows that continuations from the student distilled with the KLD reaches the lowest perplexity and match GPT-2 with nucleus sampling in 32 forward passes.
In table 1, we compare the downstream performance on the tasks of Gao et al. (2021) before and after distillation. We observe that SDTT minimally affects the results, and that student distilled with the KLD objective reaches higher accuracies than other students in all but one task
Figure 4(a) measures the diversity of samples using the self-BLEU score (Zhu et al., 2018), for the students distilled with the KLD objective. See table 1 for results with the MSE and TVD. We find that SDTT minimally decreases the diversity. Compared to distilling autoregressive models (Agarwal et al., 2024), SDTT minimally reduces the diversity. For reference, Agarwal et al. (2024) routinely observes an increase of 15 in self-BLEU while we observe a change of at most 2 for the KLD student. See table 1 for more results and details on the self-BLEU score.
Figure 6 shows that students distilled with KLD have higher unconditional generative perplexity than those distilled with the MSE. However, KLD is the only objective that preserves performance in the LAMBADA data set while still significantly reducing the generative perplexity compared to the teacher. Therefore, in the remainder of this work, we focus on the KLD.